# Dynamical patterns and nonreciprocal effective interactions in an active-passive mixture through exact hydrodynamic analysis

### v2

@@ Acknowledgements
We thank Tal Agranov, Martin Burger, Mike Cates, Clement Erignoux, Sarah Loos, and Johannes Zimmer for helpful discussions. M. Bruna was supported by a Royal Society University Research Fellowship (grant no. URF/R1/180040). J. Mason was supported by the Royal Society Award (RGF/EA/181043) and the Cantab Capital Institute for the Mathematics of Information of the University of Cambridge. For the numerical work, we used the Julia programming language [60] and the following packages and tools: DrWatson.jl [61], DifferentialEquations.jl [62], NonlinearSolve.jl [63]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any author accepted manuscript arising from this submission.
@@ Appendix A Hydrodynamic limit
To obtain these equations from the particle-based dynamics of the APLG, we define $\eta_{\sigma}(\bm{x},t)$ to be $1$ if there is a $\sigma$-particle at position $\bm{x}$ and 00 otherwise. The local density is formally defined as the mean number of particles in a mesoscopic box of radius $r$ around $\bm{x}$,
$\hat{\rho}_{\sigma}(\bm{x},t)\approx\frac{1}{(2r/h+1)^{2}}\sum_{\|\bm{x}-\bm{y}\|_{\infty}<r}\eta_{\sigma}(\bm{y},t),$ (S2)
@@
This is proven in the APLG by generalizing the work of Erignoux [19]. That work considered a system of pure active particles with continuously varying orientations $(\cos\theta,\sin\theta)$ with $\theta\in[0,2\pi)$ instead of only $\theta=\{0,\pi\}$ used here. In both cases, the proof of the convergence $\hat{\rho}_{\sigma}\to\rho_{\sigma}$ is technically challenging because the models are of nongradient type in the sense of [18]. It is worth emphasizing that this classification is separate from whether the model can be derived as a gradient flow of an equilibrium free energy. Indeed, the hydrodynamic limit of a mixture of two species undergoing a symmetric simple exclusion process is a gradient flow, but the underlying microscopic model is of nongradient type [26]. Being of nongradient type means instead that the current $j_{\bm{x},\bm{x}+h\bm{e}_{i}}$ from site $\bm{x}$ to $\bm{x}+h\bm{e}_{i}$, cannot be written as the discrete difference of a local function $g$
$j_{\bm{x},\bm{x}+h\bm{e}_{i}}\neq g_{\bm{x}+h\bm{e}_{i}}(\eta)-g_{\bm{x}}(\eta).$ (S3)
@@
The nongradient method [26] involves projecting the current onto a space of discrete differences and proving that it can be replaced by its local average in the hydrodynamic limit, e.g., the instantaneous current of $\sigma$-particles going between neighboring sites $\bm{x}$ and $\bm{x}+h\bm{e}_{1}$ can be approximated as
$j_{\bm{x},\bm{x}+h{\bm{e}}_{1}}^{\sigma}\simeq d_{s}(\rho)[\eta_{\sigma}(\bm{x}+h\bm{e}_{1})-\eta_{\sigma}(\bm{x})]+{\cal D}(\rho)[\eta(\bm{x}+h\bm{e}_{1})-\eta(\bm{x})]+\text{Pe}\left[\sigma d_{s}(\rho)\eta_{\sigma}(\bm{x})+s(\rho)m\eta_{\sigma}(\bm{x})\right],$ (S4)
@@
Another signature of the model being of nongradient type is that its hydrodynamic limit contains transport coefficients whose dependence on the density is not known, albeit they are characterized by a variational formula over an infinite-dimensional space [27]. This is the case for the self-diffusion coefficient $d_{s}(\rho)$ in Eq. (S1). In our analysis, we use the approximation (3) of $d_{s}(\rho)$ obtained in [29], exploiting a rigorous recursive approach proposed in [64]. The idea equates to performing a Taylor expansion of $d_{s}(\rho)$ around $\rho=0,1$ up to the first order and combining the two linear approximations in the minimal cubic polynomial. At the level of the self-diffusion coefficient, the mean-field approximation $\langle\eta_{\sigma}(\bm{x},t)\eta_{\sigma}(\hat{\bm{x}},t)\rangle\approx\rho_{\sigma}(\bm{x},t)\rho_{\sigma}(\hat{\bm{x}},t)$ would result in $d_{s}(\rho)\equiv 1-\rho$ in (3) and, in turn, $\mathcal{D}(\rho)\equiv 1$ and $s(\rho)\equiv 0$ in (2).
@@ 1 Ordinary differential equations
where $\Gamma$ is the friction matrix, which is real, symmetric, and positive definite. The energy $E$ is nonincreasing under the dynamics (S6). Assume that $\textbf{x}=0$ is a fixed point [$\nabla E(0)=0$] and linearise (S6):
@@
where $H_{0}$ is the Hessian of $E$ at $\textbf{x}=0$ and $\Gamma_{0}=\Gamma(0)$. Stability depends on the eigenvalues of $B$. Using that $\Gamma_{0}$ is real symmetric positive definite and similarity-transforming with $\Gamma_{0}^{1/2}$, the eigenvalues of $B$ are those of the real symmetric matrix $-\Gamma_{0}^{-1/2}H_{0}\Gamma_{0}^{-1/2}$, hence they are always real. If the eigenvalues of $H_{0}$ are positive (minimum of $E$), then the eigenvalues of $B$ are negative (stable fixed point, as expected). The translation invariance (S5) means that $H_{0}$ has a zero eigenvalue, but this may always by removed by defining positions relative to the center of mass.
@@
$\dot{\textbf{x}}=\Gamma^{-1}\textbf{F},$ (S10)
where F is a force, which is not generically the gradient of any energy $E$. We continue to assume that $\textbf{x}=0$ is a fixed point and translation invariance (now in the form $\sum_{j}(\partial F_{i}/\partial x_{j})=0$ for all $i$). Linearising about the fixed point gives the analog of (S8)
@@
Consider three particles with positions $x_{1},x_{2},x_{3}$ moving in the interval $[0,2\pi)$ with periodic boundaries. We take as the equation of motion
@@ 2 Cross-diffusion systems
where ${\cal B}={\cal B}(\rho_{1},\rho_{2})$ is a matrix. It is not obvious in general whether such an equation can be factorised into the reciprocal form (S17). However, if the eigenvalues of ${\cal B}$ are complex the system can not be factorised in this way.
Physically, the most common signature of nonreciprocal effective interactions is that species 1 is attracted to species 2, while species 2 is repelled from species 1. This intuitively corresponds to ${\cal B}_{12}$ and ${\cal B}_{21}$ having opposite signs. Note however that this difference of signs may also occur in reciprocal systems, [^1] For example, $MH=\begin{pmatrix}8&3\\ 3&8\end{pmatrix}\begin{pmatrix}4&-1\\ -1&2\end{pmatrix}=\begin{pmatrix}29&-2\\ 4&13\end{pmatrix}={\cal B}$. so it cannot be regarded as a definite signature of nonreciprocity.
@@
$\partial_{t}\begin{pmatrix}\rho_{1}\\ \rho_{2}\end{pmatrix}=-\nabla\cdot\left[M\begin{pmatrix}\bm{F}_{1}\\ \bm{F}_{2}\end{pmatrix}\right],\qquad\begin{pmatrix}\bm{F}_{1}\\ \bm{F}_{2}\end{pmatrix}=-G\begin{pmatrix}\nabla\rho_{1}\\ \nabla\rho_{2}\end{pmatrix},$ (S20)
where $G$ is a nonsymmetric matrix but $M$ is still symmetric positive definite, and $\bm{F}_{1},\bm{F}_{2}$ are the “forces” on species 1, 2: If $G_{12}<0$ and $G_{21}>0$ then species 1 feels a force towards 2, while species 2 feels a force away from 1. The systems that we consider reduce to equilibrium (reciprocal) systems in some suitable limit in which $M$ is known. We suppose the nonreciprocity enters via the forces $G$ so the matrix $M$ is fixed to its equilibrium form. Then:
@@
$\partial_{t}\rho_{a}&=\partial_{x}\left[(d_{s}+\rho_{a}\mathcal{D})\partial_{x}\rho_{a}+\rho_{a}\mathcal{D}\partial_{x}\rho_{0}+\frac{1}{2}{\rm Pe}^{2}(s\rho_{a}+d_{s})\partial_{x}(d_{s}\rho_{a})\right],\\ \partial_{t}\rho_{0}&=\partial_{x}\left[(d_{s}+\rho_{0}\mathcal{D})\partial_{x}\rho_{0}+\rho_{0}\mathcal{D}\partial_{x}\rho_{a}+\frac{1}{2}\text{Pe}^{2}s\rho_{0}\partial_{x}(d_{s}\rho_{a})\right],\end{aligned}$ (S24)
@@
Note that, in this case, $MH$ is not symmetric but the system is reciprocal (since $H$ is symmetric).
@@
Figure S1: Illustrative features of the spectrum of $MG=MH+\frac{1}{2}\text{Pe}^{2}MG_{\text{NR}}$ as defined in (S27), (S28), for various values of Pe ($\text{Pe}=2.5,5,7.5$). Denoting the eigenvalues of $MG$ by $\lambda_{1},\lambda_{2}$, the blue region indicates complex eigenvalues (with discriminant $\Delta=(\lambda_{1}+\lambda_{2})^{2}-4\lambda_{1}\lambda_{2}<0$) and the orange region indicates that at least one eigenvalue has negative real part (Re$(\lambda_{i})<0$). Where the blue and orange regions overlap, the instability involves traveling and growing sinusoids; otherwise, the orange region corresponds to sinusoids that grow but do not travel.
@@
Appendix C The method of coexisting phases and the binodal curve
@@
[…]
@@
Appendix D Linear stability of homogeneous solutions and the spinodal curve
@@
$\lambda A_{1}&=-q^{2}A_{1}-iq\text{Pe}(1-\phi)A_{3},\\ \lambda A_{2}&=-q^{2}\phi_{a}{\cal D}(\phi)A_{1}-q^{2}d_{s}(\phi)A_{2}-iq\text{Pe}[\phi_{a}s(\phi)+d_{s}(\phi)]A_{3},\\ \lambda A_{3}&=-iq\text{Pe}d_{s}^{\prime}(\phi)\phi_{a}A_{1}-iq\text{Pe}d_{s}(\phi)A_{2}-[q^{2}d_{s}(\phi)+2]A_{3}.\end{aligned}$ (S46)
@@
Figure S2: Eigenvalues of $W$ (S47) at the onset of instability. Active volume fraction $\phi_{a}$ increases from left to right. The critical volume fraction is displayed in the central column. (a) Cahn-Hiliard instability (stationary, large-scale). (b) Conserved-Hopf instability (oscillatory, large-scale). (c) Conserved-wave instability (oscillatory, small-scale).
@@ 1 Spinodal curves
Figure S3: Phase diagrams spanned by $\phi$ and $\nu$. The Spinodal (Solid Blue) encloses the region of linear stability of homogeneous solutions. The dashed (solid) blue line indicates the minimum (maximum) of (S50) and (S51). The Binodal (Red) encloses the region of Phase separation. (a-b) spinodal is contained within the binodal (c-f) spinodal protrudes through the binodal, creating instability in PS solutions.
@@ 3 Small $q$ approximation of the spinodal
where $\widetilde{W}=-q^{2}M(H+\frac{\text{Pe}^{2}}{2}G_{\text{NR}})$ and $M,H,G_{\text{NR}}$ coincide with the matrices in (S27,S28) corresponding to the nonreciprocal form of the outer problem. Hence, the matrix of linear stability in the limit $q\ll 1$ coincides, up to the factor $-q^{2}$, with the matrix product $MG$ when writing the outer problem as a cross-diffusion problem (Example 3).
@@
Appendix E Numerical methods for the hydrodynamic PDEs
@@ 2 Traveling solutions in the finite system
So far, we have $3N$ equations ($3(N-1)$ in (S66) and three in (S67)) for $13N+$ 1 unknowns. The remaining degree of freedom is removed by noting the problem has translational symmetry; hence, we set $\varrho_{0,0}=\phi_{p}$ without loss of generality.
The nonlinear system of $3N$ equations (S66)-(S67) is then solved numerically using NonlinearSolve() from the Julia NonlinearSolve.jl package [63], with parameters reltol=1e-8, abstol=1e-8 and maxiters=20. The solver first tries less robust Quasi-Newton methods for more performance and then tries more robust techniques if the faster ones fail. As such, it requires a close enough initial guess for convergence. The first time, we initialize the iterative solver with the long-time ($T=1000$) solution of the time-dependent problem (Subsec. 1) with parameters $\phi=0.67,\phi_{a}=0.37$,$\text{Pe}=7.5$, $L=25$,$N=500$. Once we have found a traveling solution to (S66,S67), we use it as an initial condition for the problem with slightly altered parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$, as well as larger domains, e.g., $L^{\prime}=2L$ or $N^{\prime}=2N$.

### v3

@@ Acknowledgements
We thank Tal Agranov, Martin Burger, Mike Cates, Clement Erignoux, Sarah Loos, and Johannes Zimmer for helpful discussions. M. Bruna was supported by a Royal Society University Research Fellowship (grant no. URF/R1/180040). J. Mason was supported by the Royal Society Award (RGF/EA/181043) and the Cantab Capital Institute for the Mathematics of Information of the University of Cambridge. For the numerical work, we used the Julia programming language [60] and the following packages and tools: DrWatson.jl [61], DifferentialEquations.jl [62], NonlinearSolve.jl [63].
@@ A Hydrodynamic limit
To obtain these equations from the particle-based dynamics of the APLG, we define $\eta_{\sigma}({\mathbf{x}},t)$ to be $1$ if there is a $\sigma$-particle at position $\mathbf{x}$} and 0 otherwise. The local density is formally defined as the mean number of particles in a mesoscopic box of radius $r$ around $\mathbf{x}$},
$\hat{\rho}_{\sigma}({\mathbf{x}},t)\approx\frac{1}{(2r/h+1)^{2}}\sum_{\|{\mathbf{x}}-{\mathbf{y}}\|_{\infty}<r}\eta_{\sigma}({\mathbf{y}},t),$ (S2)
@@
This is proven in the APLG by generalizing the work of Erignoux [19]. That work considered a system of pure active particles with continuously varying orientations $(\cos\theta,\sin\theta)$ with $\theta\in[0,2\pi)$ instead of only $\theta=\{0,\pi\}$ used here. In both cases, the proof of the convergence $\hat{\rho}_{\sigma}\to\rho_{\sigma}$ is technically challenging because the models are of nongradient type in the sense of [18]. It is worth emphasizing that this classification is separate from whether the model can be derived as a gradient flow of an equilibrium free energy. Indeed, the hydrodynamic limit of a mixture of two species undergoing a symmetric simple exclusion process is a gradient flow, but the underlying microscopic model is of nongradient type [26]. Being of nongradient type means instead that the current $j_{{\mathbf{x}},{\mathbf{x}}+h{\mathbf{e}}_{i}}$ from site $\mathbf{x}$} to ${\mathbf{x}}+h{\mathbf{e}}_{i}$, cannot be written as the discrete difference of a local function $g$
$j_{{\mathbf{x}},{\mathbf{x}}+h{\mathbf{e}}_{i}}\neq g_{{\mathbf{x}}+h{\mathbf{e}}_{i}}(\eta)-g_{{\mathbf{x}}}(\eta).$ (S3)
@@
The nongradient method [26] involves projecting the current onto a space of discrete differences and proving that it can be replaced by its local average in the hydrodynamic limit, e.g., the instantaneous current of $\sigma$-particles going between neighboring sites $\mathbf{x}$} and ${\mathbf{x}}+h{\mathbf{e}}_{1}$ can be approximated as
$j_{{\mathbf{x}},{\mathbf{x}}+h{{\mathbf{e}}}_{1}}^{\sigma}\simeq d_{s}(\rho)[\eta_{\sigma}({\mathbf{x}}+h{\mathbf{e}}_{1})-\eta_{\sigma}({\mathbf{x}})]+{\cal D}(\rho)[\eta({\mathbf{x}}+h{\mathbf{e}}_{1})-\eta({\mathbf{x}})]+\text{Pe}\left[\sigma d_{s}(\rho)\eta_{\sigma}({\mathbf{x}})+s(\rho)m\eta_{\sigma}({\mathbf{x}})\right],$ (S4)
@@
Another signature of the model being of nongradient type is that its hydrodynamic limit contains transport coefficients whose dependence on the density is not known, albeit they are characterized by a variational formula over an infinite-dimensional space [27]. This is the case for the self-diffusion coefficient $d_{s}(\rho)$ in Eq. (S1). In our analysis, we use the approximation (3) of $d_{s}(\rho)$ obtained in [29], exploiting a rigorous recursive approach proposed in [64]. The idea equates to performing a Taylor expansion of $d_{s}(\rho)$ around $\rho=0,1$ up to the first order and combining the two linear approximations in the minimal cubic polynomial. At the level of the self-diffusion coefficient, the mean-field approximation $\langle\eta_{\sigma}({\mathbf{x}},t)\eta_{\sigma}(\hat{\mathbf{x}},t)\rangle\approx\rho_{\sigma}({\mathbf{x}},t)\rho_{\sigma}(\hat{\mathbf{x}},t)$ would result in $d_{s}(\rho)\equiv 1-\rho$ in (3) and, in turn, $\mathcal{D}(\rho)\equiv 1$ and $s(\rho)\equiv 0$ in (2).
@@ 1 Ordinary differential equations
where $\Gamma$ is the friction matrix, which is real, symmetric, and positive definite. The energy $E$ is nonincreasing under the dynamics (S6). Assume that $\textbf{x}=0$ is a fixed point $\nabla E(0)=0$] and linearise (S6):
@@
where $H_{0}$ is the Hessian of $E$ at $\textbf{x}=0$ and $\Gamma_{0}=\Gamma(0)$. Stability depends on the eigenvalues of $B$. Using that $\Gamma_{0}$ is real symmetric positive definite and similarity-transforming with $\Gamma_{0}^{1/2}$, the eigenvalues of $B$ are those of the real symmetric matrix $\Gamma_{0}^{-1/2}H_{0}\Gamma_{0}^{-1/2}$, hence they are always real. If the eigenvalues of $H_{0}$ are positive (minimum of $E$), then the eigenvalues of $B$ are negative (stable fixed point, as expected). The translation invariance (S5) means that $H_{0}$ has a zero eigenvalue, but this may always by removed by defining positions relative to the center of mass.
@@
$\dot{\textbf{x}}=\Gamma^{-1}{\mathbf{F}},$ (S10)
where $\mathbf{F}$} is a force, which is not generically the gradient of any energy $E$. We continue to assume that $\textbf{x}=0$ is a fixed point and translation invariance (now in the form $\sum_{j}(\partial F_{i}/\partial x_{j})=0$ for all $i$). Linearising about the fixed point gives the analog of (S8)
@@
Consider three particles with positions $x_{1},x_{2},x_{3}$ moving in the interval $0,2\pi)$ with periodic boundaries. We take as the equation of motion
@@ 2 Cross-diffusion systems
[…]
@@
$\partial_{t}\begin{pmatrix}\rho_{1}\\ \rho_{2}\end{pmatrix}=-\nabla\cdot\left[M\begin{pmatrix}{\mathbf{F}}_{1}\\ {\mathbf{F}}_{2}\end{pmatrix}\right],\qquad\begin{pmatrix}{\mathbf{F}}_{1}\\ {\mathbf{F}}_{2}\end{pmatrix}=-G\begin{pmatrix}\nabla\rho_{1}\\ \nabla\rho_{2}\end{pmatrix}$, (S20)
where $G$ is a nonsymmetric matrix but $M$ is still symmetric positive definite, and ${\mathbf{F}}_{1},{\mathbf{F}}_{2}$ are the “forces” on species 1, 2: If $G_{12}<0$ and $G_{21}>0$ then species 1 feels a force towards 2, while species 2 feels a force away from 1. The systems that we consider reduce to equilibrium (reciprocal) systems in some suitable limit in which $M$ is known. We suppose the nonreciprocity enters via the forces $G$ so the matrix $M$ is fixed to its equilibrium form. Then:
@@
$\begin{aligned} \partial_{t}\rho_{a}&=\partial_{x}\left[(d_{s}+\rho_{a}\mathcal{D})\partial_{x}\rho_{a}+\rho_{a}\mathcal{D}\partial_{x}\rho_{0}+\frac{1}{2}{\rm Pe}^{2}(s\rho_{a}+d_{s})\partial_{x}(d_{s}\rho_{a})\right],\\ \partial_{t}\rho_{0}&=\partial_{x}\left[(d_{s}+\rho_{0}\mathcal{D})\partial_{x}\rho_{0}+\rho_{0}\mathcal{D}\partial_{x}\rho_{a}+\frac{1}{2}\text{Pe}^{2}s\rho_{0}\partial_{x}(d_{s}\rho_{a})\right],\end{aligned}$ (S24)
@@
Note that, in this case, MHMH is not symmetric but the system is reciprocal (since $H$ is symmetric).
@@
[…]
@@
C The method of coexisting phases and the binodal curve
@@
$\begin{aligned} \Phi(R(\phi_{l}))-\Phi(R(\phi_{v}))&=\int_{x_{v}}^{x_{l}}\Phi^{\prime}(R(\rho))R^{\prime}(\rho)\partial_{x}\rho\,dx=\int_{x_{v}}^{x_{l}}g_{0}(\rho)R^{\prime}(\rho)\partial_{x}\rho\,dx\\ &=\int_{x_{v}}^{x_{l}}g(\rho,\partial_{x}\rho,\partial_{x}^{2}\rho)R^{\prime}(\rho)\partial_{x}\rho\,dx+\int_{x_{v}}^{x_{l}}[\kappa(\rho)\partial_{x}^{2}\rho-\Lambda(\rho)(\partial_{x}\rho)^{2}]R^{\prime}(\rho)\partial_{x}\rho\,\mathrm{d}x,\end{aligned}$ (S36)
@@
D Linear stability of homogeneous solutions and the spinodal curve
@@
$\begin{aligned} \lambda A_{1}&=-q^{2}A_{1}-iq\text{Pe}(1-\phi)A_{3},\\ \lambda A_{2}&=-q^{2}\phi_{a}{\cal D}(\phi)A_{1}-q^{2}d_{s}(\phi)A_{2}-iq\text{Pe}[\phi_{a}s(\phi)+d_{s}(\phi)]A_{3},\\ \lambda A_{3}&=-iq\text{Pe}d_{s}^{\prime}(\phi)\phi_{a}A_{1}-iq\text{Pe}d_{s}(\phi)A_{2}-[q^{2}d_{s}(\phi)+2]A_{3}.\end{aligned}$ (S46)
@@
Figure S2: Linear stability of the APLG system. Eigenvalues of $W$ (S47) at the onset of instability. Active volume fraction $\phi_{a}$ increases from left to right. The critical volume fraction is displayed in the central column. (a) Cahn-Hiliard instability (stationary, large-scale). (b) Conserved-Hopf instability (oscillatory, large-scale). (c) Conserved-wave instability (oscillatory, small-scale).
@@ 1 Spinodal curves
Figure S3: Protrusion of the spinodal beyond the binodal. Phase diagrams spanned by $\phi$ and $\nu$. The spinodal (solid blue) encloses the region of linear stability of homogeneous solutions. The dashed (solid) blue line indicates the minimum (maximum) of (S50) and (S51). The binodal (red) encloses the region of phase separation. (a-b) spinodal is contained within the binodal (c-f) spinodal protrudes through the binodal, creating instability in PS solutions.
@@ 3 Small $q$ approximation of the spinodal
where $\widetilde{W}=-q^{2}M(H+\frac{\text{Pe}^{2}}{2}G_{\text{NR}})$ and $M,H,G_{\text{NR}}$ coincide with the matrices in (S27,S28) corresponding to the nonreciprocal form of the outer problem. Hence, the matrix of linear stability in the limit $q\ll 1$ coincides, up to the factor $q^{2}$, with the matrix product $MG$ when writing the outer problem as a cross-diffusion problem (Example 3).
@@
E Numerical methods for the hydrodynamic PDEs
@@ 2 Traveling solutions in the finite system
So far, we have $3N$ equations (3(N1)3(N-1) in (S66) and three in (S67)) for $3N+1$ unknowns. The remaining degree of freedom is removed by noting the problem has translational symmetry; hence, we set $\varrho_{0,0}=\phi_{p}$ without loss of generality.
The nonlinear system of $3N$ equations (S66)-(S67) is then solved numerically using NonlinearSolve() from the Julia NonlinearSolve.jl package [63], with parameters reltol=1e-8, abstol=1e-8 and maxiters=20. The solver first tries less robust Quasi-Newton methods for more performance and then tries more robust techniques if the faster ones fail. As such, it requires a close enough initial guess for convergence. The first time, we initialize the iterative solver with the long-time ($T=1000$) solution of the time-dependent problem (Subsec. 1) with parameters $\phi=0.67,\phi_{a}=0.37$$\text{Pe}=7.5$, $L=25$,$N=500$. Once we have found a traveling solution to (S66,S67), we use it as an initial condition for the problem with slightly altered parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$, as well as larger domains, e.g., $L^{\prime}=2L$ or $N^{\prime}=2N$.